Maxis taking a cross-country road trip. Gas prices vary as the friends travel across the US from $4 dollars per gallon on the east coast, to $3 in the mid-US, to $5 on the west coast. If they used twice as much gas in the mid-US than on either coasts combined, and they spend $515 on gas to purchased 150 gallons of gas, how many gallons of gas did they buy at each price?
Answer by Boreal(15235)

x=gallons of gas in midwest
150-x on both coasts
x=2(150-x)=300-2x
3x=300
x=100 gallons in midwest or $300 ANSWER
That leaves 50 gallons in both coasts
w=west
e=east=50-w
5w+4(50-w)=215, 515-300.
5w+200-4w=215
w=15 gallons west coast ($75) ANSWER
e=35 gallons ($140) ANSWER